# Impact of the Pandemic on Currency Circulation in Brazil: Projections using the SARIMA Model

### v1

@@ 1.1.2 The economic impact of the pandemic in Brazil
    Economic stimulus policies: To mitigate the economic effects, the federal government implemented emergency measures such as the Emergency Aid program, which injected billions of reais directly into the economy, increasing currency circulation. This program was crucial in sustaining the consumption of low-income families during the crisis;
@@
    Changes in consumption patterns and money usage: During the pandemic, there was a rise in demand for physical cash, possibly due to economic uncertainty and concerns about access to banking services. Additionally, the crisis accelerated financial digitalization, with a higher use of electronic payments and instant transfers, such as Pix.
@@ 3.1 Currency Circulation (2000–2023)
  • Test set: Includes the subsequent 12 months (from April 2020 to March 2021), used to validate the model by evaluating its performance with metrics such as MAPE (Mean Absolute Percentage Error) and RMSE (Root Mean Squared Error);
@@ A.2 Estimated Coefficients and Standard Errors
Table 4 presents the estimated coefficients for the non-seasonal and seasonal components of the model, along with their corresponding standard errors.
Table 4: Estimated Coefficients and Standard Errors of the ARIMA Model
@@
Seasonal Part
@@ A.3 Model Performance Measures
Table 5: Model Performance Measures of the ARIMA Model

### v2

@@ 1.1.2 The economic impact of the pandemic in Brazil
    Economic stimulus policies: To mitigate the economic effects, the federal government implemented emergency measures such as the Emergency Aid program, which injected billions of reais directly into the economy, increasing currency circulation. This program was crucial in sustaining the consumption of low-income families during the crisis [15];
@@
    Changes in consumption patterns and money usage: During the pandemic, there was a rise in demand for physical cash, possibly due to economic uncertainty and concerns about access to banking services. Additionally, the crisis accelerated financial digitalization, with a higher use of electronic payments and instant transfers, such as Pix [13] [14].
@@ 3.1 Currency Circulation (2000–2023)
  • Test set: Includes the subsequent 12 months (from April 2020 to March 2021), used to validate the model by evaluating its performance with metrics such as MSE (Mean Squared Error), MAPE (Mean Absolute Percentage Error), and RMSE (Root Mean Squared Error);
@@ A.2 Estimated Coefficients and Standard Errors
Tables 4 and 5 present the estimated coefficients for the non-seasonal and seasonal components of the model, along with their corresponding standard errors.
Table 4: Estimated Coefficients and Standard Errors of the SARIMA Model (Non-seasonal Part)
@@
Table 5: Estimated Coefficients and Standard Errors of the SARIMA Model (Seasonal Part)
Parameter Estimated Value Standard Error
@@ A.3 Model Performance Measures
Table 6: Model Performance Measures of the SARIMA Model